"five more than a number is the same as 4 times the sum of the number and 8"
I dont know why, but i can get it! it must be than dang "more than'...
Thank you for your time!

Answer by Alan3354(69443)

"five more than a number is the same as 4 times the sum of the number and 8"
----
n = a number
n + 5 = 4(n + 8)
n + 5 = 4n + 32
3n = -27
n = -9